Isabel is 7 years older than Marie in 6 years, Isabel’s age will be 11 years less than twice Marie’s age. How old is each now?
Answer by ikleyn(52781)

.

            Hey, the punctuation signs were invented/introduced to use them properly (!)


Marie's age = x

Isabel's age = (x+7)


In 6 years

    (x+7) + 6 = 2*(x+6) - 11

    
     x + 13 = 2x + 12 - 11

     x + 13 = 2x + 1

     13 - 1 = 2x - x

     x       = 12.


ANSWER.  Marie is 12 years old;  Isabel is  12+7 = 19 years old.

Solved.

----------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.